OpenStudy (anonymous):

A clock has a 14 ft long minute hand. What is the angular speed and linear speed of its tip? How far does its tip move in half an hour?

OpenStudy (anonymous):

@ganeshie8 could you give me a hand with this? I have a final exam tomorrow so I"m trying to review everything

OpenStudy (anonymous):

do you know the perimeter of a circle?

OpenStudy (anonymous):

2*pi*r

OpenStudy (anonymous):

r = radius = 2(pi)(14) all devided by half because its only half a circle so you would get 14pi

OpenStudy (anonymous):

i don't know what angular vs linear speed is, so maybe that will get you started

ganeshie8 (ganeshie8):

\[\text{Anglular speed } = \dfrac{2 \pi}{T} = \dfrac{2\pi }{\text{1 hour}} = 2\pi \text{ radians/hour}\]

ganeshie8 (ganeshie8):

\[\text{Linear speed } = \dfrac{2 \pi ~r}{T} = \dfrac{2\pi\times 14 }{\text{1 hour}} = 28 \pi \text{ ft/hour}\]